Write Verilog codes of 8-bit Log Left Shifter and 4-bit Barrel Right/Left Shifter.

Simulate the designs with test benches of at least 4 test cases. 


1
Expert's answer
2021-09-15T02:13:35-0400

A barrel shifter is simply a bit-rotating shift register.

The bits shifted out of LSB end of the register are

shifted back into the MSB end of the register. The

number of multiplexers required is n*log2(n), for an n

bit word. Four common word sizes and the number of

multiplexers needed are listed below:

64_bit -64*log2(64)=64*6=364

32_bit -32*log2(32)=32*5=160

16_bit -16*log2(16)=16*4=64

8_bit -8*log2(8)=64*3=24